OpenStudy (yacoub1993):

Find the area of a sector with a central angle of 170° and a diameter of 9.1 cm. Round to the nearest tenth.

OpenStudy (anonymous):

diameter=9.1 cm radius =9.1/2=4.55 cm required area \[=\pi r ^{2}*\frac{ 170 }{ 360 }\]

OpenStudy (anonymous):

u are asking direct formula based questions... read theory and formulas and den u will solve dem directly.. area= (angle*r^2)/2 where angle is in radian

OpenStudy (yacoub1993):

@gauravkj i got 1759.7125

OpenStudy (anonymous):

angle is in radian... 170/180 is the value of angle

OpenStudy (yacoub1993):

@gauravkj i didnt get it

OpenStudy (yacoub1993):

u der

OpenStudy (anonymous):

which part u did not get?

OpenStudy (yacoub1993):

area= (angle*r^2)/2 .... (170*4.55^2)/2

OpenStudy (yacoub1993):

am i right

OpenStudy (anonymous):

angle is not equals to 170.. it is equal to 170/180; see it like this.. area of a circle is pie*r*r.. and it is 360 degrees.. urs is 170 degrees.. so area should be = (area of circle/360)*170

OpenStudy (yacoub1993):

then

OpenStudy (anonymous):

then area= (170/180)*(radius^2)/2

OpenStudy (yacoub1993):

9.776 i got

OpenStudy (yacoub1993):

lol this is soooooo hard

OpenStudy (yacoub1993):

A. 122.9 cm^2 B. 30.7 cm^2 C. 8.6 cm^2 D. 3.4 cm^2

OpenStudy (anonymous):

9.77 cm^2 O.o u sure these are the only options? lol

OpenStudy (yacoub1993):

yess i am sure i will snap a pic and show u

OpenStudy (yacoub1993):

http://prntscr.com/2s5b61

OpenStudy (anonymous):

ohh sorry...... angle is (170/180)*pie.. so silly of me...

OpenStudy (anonymous):

u will get 9.776*3.14 times.. so B option

OpenStudy (anonymous):

when i told u by that full circle analogy, i was correct then.. made mistake in converting degrees into radians..

OpenStudy (anonymous):

lol thanx.. sorry i could not make it easy and interesting for u..

OpenStudy (yacoub1993):

hehe its ok

OpenStudy (anonymous):

area=\[\pi r ^{2}*\frac{ 170 }{ 360 }=3.14*(4.55)^{2}*\frac{ 17 }{ 36 }=3.14*20.70*\frac{ 17 }{ 36 }\] =30.69=30.7cm^2